OpenStudy (anonymous):

if (-2,5) , ( 9,-4) is connected what is the slope of the line ?

OpenStudy (ahsome):

Equation you would use: \[m=\frac{y^1-y^2}{x^1-x^2}\] \[(-2,5)=(x^1,y^1)\] \[(9,-4)=(x^2,y^2)\] Sub in those values to the equation to get \(m\)

OpenStudy (ahsome):

Does that make sense @UnderCover22 ?

OpenStudy (anonymous):

nope it dosent

OpenStudy (anonymous):

No not really

OpenStudy (ahsome):

Um...

OpenStudy (ahsome):

Ok, do you know that the slope of a line is \(m\)?

OpenStudy (anonymous):

my answer ended up -11 over 1 and i think i did something wrong

OpenStudy (anonymous):

@undercover please firs plot the graph and then choose a point joining the two lines and then look for the slope

OpenStudy (anonymous):

i did now i need to know what is the slope of the line ?

OpenStudy (anonymous):

The line is diagonal

OpenStudy (anonymous):